# GARG-AML against Smurfing: A Scalable and Interpretable Graph-Based Framework for Anti-Money Laundering

### v1

@@ 3.3 Directed
A defining trait of smurfing is the unidirectional flow of funds—money moves from one account to others with minimal or no return flow [39, 24, 27, 38], forming what we call pure smurfing patterns (see Figure 3). To capture this behaviour, we extend GARG-AML to directed networks.
Figure 3: The directed version of the smurfing pattern example, extended with an additional sending node (A’) and receiving node (E’). Nodes at level 0 are put in blue, at level 1 in purple and at level 2 in green.
In the directed network of Figure 3, unlike the undirected case (Figure 2), node $A$ only sends funds to smurfs, while node $E$ only receives them. This asymmetry leads to a non-symmetric adjacency matrix, requiring nine matrix blocks instead of three (see Equation (10)). To ensure consistency in block-wise density calculations—especially with multiple senders or receivers—we distinguish senders (level 0) from receivers (level 2), as shown in Figure 3.
Accordingly, we define two types of second-order neighbourhoods in network $G(V,E)$. The strong second-order neighbourhood, denoted $\mathcal{N}^{s}_{2}(v)$, considers nodes reachable from $v$ via a directed path of length two. The weak second-order neighbourhood, $\mathcal{N}^{w}_{2}(v)$, uses the undirected network. In both, only nodes at exact distance two are included.
@@
This captures nodes that are connected at distance two in the undirected sense, but not reachable via directed paths. To account for reverse flow patterns—such as when node $E$ in Figure 3 is the focal node—we repeat this selection on the reversed network. This ensures that corresponding nodes like $E^{\prime}$ are correctly placed at level 0 when node $E$ would be considered.
We take the sizes to be $\mid\mathcal{N}_{1}(v)\mid=n$, $\mid\mathcal{N}^{s}_{2}(v)\mid=m$, and $\mid\mathcal{N}^{w}_{2}(v)\setminus\mathcal{N}^{s}_{2}(v)\mid=l-1$, resulting in $l$ nodes at level 0, $n$ nodes at level 1, and $m$ nodes at level 2. The resulting adjacency matrix for the network on the right in Figure 3 is then
As in the undirected case, structural constraints of the adjacency matrix are considered when calculating block-wise densities. For the upper-left block, $\text{score}_{00}$, the first row and column are excluded since all other level-0 nodes are second-order neighbours of the central node. Additionally, diagonal elements in all on-diagonal blocks ($\text{score}_{00}$, $\text{score}_{11}$ and $\text{score}_{22}$) are excluded from density calculations, as self-loops are not present.
Adjustments are also made to $\text{score}_{20}$ and $\text{score}_{02}$, which represent connections between the central node and second-order neighbours. The first row of $\text{block}_{20}$ and first column of $\text{block}_{02}$ are always zero and are therefore excluded.
@@
The full adjacency matrix is thus partitioned into nine blocks, as shown in Equation (11), with block labels using zero-based indexing to reflect node levels 0, 1, and 2.
@@
To compute the final score, we take the weighted average of the two blocks expected to have high density — according to the smurfing pattern in Equation (10) — and subtract the weighted average of the remaining blocks, which are expected to be sparse. As with the undirected case, a weighted average is used to account for block size, though a simple arithmetic mean is also a viable alternative. The use of the average ensures that the final GARG-AML score is between -1 and 1, to have consistent comparison and interpretation across datasets.
@@
The GARG-AML scores offer two key advantages. First, they approach 1 when a node’s second-order neighbourhood resembles a smurfing pattern, providing a clear and interpretable signal for both automated detection and human investigation. This makes them easy to integrate into existing business rules or investigative workflows. Second, the method is highly scalable. Each score is computed independently using only a node’s second-order neighbourhood, allowing for efficient parallel computation and low memory usage, even on large transaction networks.
@@ 4.1.1 Synthetic Data
We generate synthetic networks of varying sizes using three generation methods, i.e., Barabási-Albert [4], Erdős-Rényi [15] and Watts–Strogatz [44]. The resulting networks have a range of different properties, e.g., scale-free, small world, etc., that allow the evaluation of GARG-AML across different network topologies.
@@
[…]
@@ 4.1.2 Open-Source Data
As post-hoc injection of specific patterns may lead to detection bias [11] and altered topology [14], we also include data without this injection in the experiments to evaluate GARG-AML.
We use the HI-Small and LI-Large datasets, made publicly available by Altman et al. [3] and Egressy et al. [14], representing a relatively small dataset with high money laundering activity and a large dataset with low money laundering activity, respectively. Table 5 summarises the number of nodes, edges and the percentage of cases labelled as money laundering. Especially LI-Large closely resembles real-world transaction networks.
@@
The two datasets are generated using the same simulator [3, 14], which models financial transactions in a virtual world involving individuals, companies, and banks. A subset of these transactions is labelled as money laundering. Since the full transaction graph is available, the authors were able to assign detailed labels, categorizing transaction streams into eight distinct types: fan-out, fan-in, gather-scatter, scatter-gather, simple cycle, random, bipartite, and stack.
@@
[…]
@@
4.2 Model Specification
The full GARG-AML calculation pipeline consists of multiple stages, beginning with pre-processing steps that prepare the network data for analysis, followed by the computation of the GARG-AML scores. We also introduce a possible extension of the GARG-AML method that applies tree-based methods. We describe this use cases in detail below to illustrate the versatility of GARG-AML. An overview of the complete pipeline is shown in Figure 5.
@@ 4.2 Model Specification
Transaction networks often exhibit a scale-free structure [39], which are characterized by the presence of a few nodes (accounts) with disproportionately high numbers of incoming and outgoing transactions. As a result, second-order neighbourhoods can become very large and can encompass an unmanageable number of nodes for score computation.
Therefore, we first partition the network into communities using the Louvain method [5]. This method consistently shows state-of-the-art performance for community detection, and scales well to large graphs [23, 50]. Communities are identified on the basis of the undirected version of the transaction network, and only intra-community edges are retained—edges between communities are discarded.
However, the Louvain method can still produce overly large communities [17, 32]. To mitigate this, we use the implementation from NetworkX [21], which includes a resolution parameter inspired by the improvements proposed by Traag et al. [41]. In our experiments, we set this parameter to 10, which effectively encourages the algorithm to form smaller, more manageable communities, thus maintaining tractability in the GARG-AML score computation.
We benchmark GARG-AML and its extension against established state-of-the-art methods, namely, AutoAudit [24] and FlowScope [27]. For both methods, we use the implementations provided by the authors and retain their default settings.
Various small adaptations are necessary. AutoAudit returns an optimal reordering of indices in the adjacency matrix. However, this reordering affects only part of the matrix and may depend on the arbitrary labelling of nodes. To ensure robustness, we repeat the experiments three times with random initialization of the node indices and report the average results. Both methods also return an ordering of transactions resulting from the $n$ different iterations. To reflect the decreasing level of certainty, the first group gets prediction $1$, the second $\frac{n-1}{n}$ and so on until the final, $n$th, group gets score $\frac{1}{n}$. All other nodes that were not returned by the methods get score 00.
@@
1:Calculate communities using Louvain method [5]
@@ 4.3 Evaluation
While AUC-ROC is a widely used metric for binary classification, it may not be sufficient for anti-money laundering (AML) datasets, which are often highly imbalanced—typically with fewer than $0.1\%$ observations flagged as laundering money. To address this imbalance, we include AUC-PR, which has been shown to be a more reliable performance metric under situations of extreme label imbalance [9, 33]. In general terms, the AUC-PR is the expected average precision when a random recall threshold is selected. In other words, the AUC-PR measures the model’s ability to identify positive examples correctly while minimising false positives across all possible threshold values. The AUC-PR specifically focuses on the model’s precision in identifying the positive class at various recall thresholds.
Scalability is assessed in terms of calculation time. In addition to runtime, memory usage is also a key factor. We also report any instances where methods fail due to memory limitations. All the experiments were conducted on the Flemish Supercomputer (VSC) with an Intel Xeon Platinum 8360Y CPU and 200 GB of memory. Each experiment was constrained to a time budget of 16 hours, the maximum on the VSC.
@@ 4.4 Statistical Test
The statistical significance of the difference in ranks is evaluated using the Friedman test [18, 19] and the post hoc Nemenyi test [31], as described by Demšar [10]. It compares the performance of the $k$ different methods across the $N$ datasets. If we take $r_{i}^{j}$ to be the rank of method $j$ on dataset $i$ for some performance metric, then the average performance across all datasets is equal to $R_{j}=\frac{1}{N}\sum_{i}r_{i}^{j}$. In our experiments, the Friedman and Nemenyi statistical tests compare the ranks of the 8 $(k)$ methods across the 66 $(N)$ synthetic datasets. We apply a significance level of $0.05$ for these tests.
@@
where critical levels $q_{\alpha}$ are provided by Demšar [10]. We visualize the results using CD diagrams [10] for the AUC-ROC and AUC-PR.

### v2

@@ 3.3 Directed
A defining trait of smurfing is the uni-directional flow of funds. Money moves from one account to others with minimal or no return flow (Starnini et al., 2021; Lee et al., 2020; Li et al., 2020; Shadrooh & Nørvåg, 2024), forming what we call pure smurfing patterns (see Figure 3). To capture this behaviour, we extend GARG-AML to directed networks.
Figure 3: The directed version of the smurfing pattern example, extended with an additional sending node ($A^{\prime}$) and receiving node ($E^{\prime}$). Nodes at level 0 are put in blue, at level 1 in purple and at level 2 in green.
We, again, start with an example. In the directed network of Figure 3, unlike the undirected case (Figure 2), node $A$ only sends funds to the smurfs, while node $E$ only receives from them. This leads to a non-symmetric adjacency matrix, requiring the analysis of nine matrix blocks instead of three (see Equation (9)). To ensure consistency in block-wise density calculations—especially with multiple senders or receivers—nodes at distance two are classified as senders (level 0) and receivers (level 2). When considering the network on the right in Figure 3, we want that node $A^{\prime}$ is put at the same level as node $A$ and node $E^{\prime}$ at the same level as node $E$. The resulting adjacency matrix is then
Accordingly, we define two types of second-order neighbourhoods in network $G(V,E)$ for the general GARG-AML score calculations. The strong second-order neighbourhood, denoted $\mathcal{N}^{s}_{2}(v)$, considers nodes reachable from $v$ via a directed path of length two. The weak second-order neighbourhood, $\mathcal{N}^{w}_{2}(v)$, is based on the undirected network. In both, only nodes at exact distance two are included.
@@
To account for reverse flow patterns we repeat this selection on the reversed network. This means that a node $w$ a distance two of node $v$ is put at level 0 if there is no directed path of length two from $v$ to $w$ nor from $w$ to $v$.
As in the undirected case, structural constraints of the adjacency matrix are considered when calculating block-wise densities. For the upper-left block, $\text{score}_{00}$, the first row and column are excluded since all other level-0 nodes are nodes at distance two. Additionally, diagonal elements in all on-diagonal blocks ($\text{score}_{00}$, $\text{score}_{11}$ and $\text{score}_{22}$) are excluded from density calculations, as self-loops are not present.
Adjustments are also made to $\text{score}_{20}$ and $\text{score}_{02}$, which represent connections between the node and its second-order neighbours. The first row of $\text{block}_{02}$ (from the node to its second-order neighbours) and first column of $\text{block}_{20}$ (from the second-order neighbours to the node) are always zero and are therefore excluded.
@@
We take the sizes to be $\mid\mathcal{N}_{1}(v)\mid=n$, $\mid\mathcal{N}^{s}_{2}(v)\mid=m$, and $\mid\mathcal{N}^{w}_{2}(v)\setminus\mathcal{N}^{s}_{2}(v)\mid=l-1$, resulting in $l$ nodes at level 0, $n$ nodes at level 1, and $m$ nodes at level 2. The full adjacency matrix is partitioned into nine blocks, as shown in Equation (11), with block labels using zero-based indexing to reflect node levels 0, 1, and 2.
@@
To compute the final score, we take the weighted average of the two blocks expected to have high density — according to the smurfing pattern in Equation (9) — and subtract the weighted average of the remaining blocks, which are expected to be sparse. The use of the average ensures that the final GARG-AML score is between -1 and 1, to have consistent comparison and interpretation across datasets.
@@
To illustrate how these calculations are done, a small example is provided in Appendix A.2.
The GARG-AML scores offer three key advantages. First, the scores are close to 1 when a node’s second-order neighbourhood resembles a smurfing pattern, providing a clear and interpretable signal for both automated detection and human investigation. Second, the scores can easy be integrated into existing workflows. The densities and final score can be seen as additional feature values that can be added when training a model or setting up business rules. Third, the method is highly scalable. Each score is computed independently using only a node’s second-order neighbourhood, allowing for efficient parallel computation and low memory usage, even on large transaction networks.
@@ 4.1.1 Synthetic Data
We generate synthetic networks of varying sizes using three generation methods, i.e., Barabási-Albert (Barabási & Albert, 1999), Erdős-Rényi (Erdős & Rényi, 1959) and Watts–Strogatz (Watts & Strogatz, 1998). The resulting networks have a range of different properties, e.g., scale-free, small world, etc., that allow the evaluation of GARG-AML across different network topologies.
@@
Barabási-Albert 100 1 - 3 145 199 46.9 $\%$
Barabási-Albert 100 2 - 3 155 345 60.0 $\%$
Barabási-Albert 100 5 - 3 141 582 50.35 $\%$
Erdős-Rényi 100 - 0.001 3 151 122 52.32 $\%$
Erdős-Rényi 100 - 0.01 3 153 155 60.78 $\%$
Watts-Strogatz 100 1 0.001 3 144 204 48.61 $\%$
Watts-Strogatz 100 1 0.01 3 146 216 52.05 $\%$
Watts-Strogatz 100 2 0.001 3 131 268 40.46 $\%$
Watts-Strogatz 100 2 0.01 3 146 316 52.74 $\%$
Watts-Strogatz 100 5 0.001 3 151 618 52.32 $\%$
Watts-Strogatz 100 5 0.01 3 145 616 53.1 $\%$
Barabási-Albert 100 1 - 5 180 291 70.0 $\%$
Barabási-Albert 100 2 - 5 179 385 69.83 $\%$
Barabási-Albert 100 5 - 5 175 678 74.29 $\%$
Erdős-Rényi 100 - 0.001 5 182 196 70.33 $\%$
Erdős-Rényi 100 - 0.01 5 163 161 65.03 $\%$
Watts-Strogatz 100 1 0.001 5 176 271 65.91 $\%$
Watts-Strogatz 100 1 0.01 5 152 248 69.74 $\%$
Watts-Strogatz 100 2 0.001 5 171 376 70.18 $\%$
Watts-Strogatz 100 2 0.01 5 180 380 67.22 $\%$
Watts-Strogatz 100 5 0.001 5 171 669 69.01 $\%$
Watts-Strogatz 100 5 0.01 5 165 655 66.67 $\%$
Barabási-Albert 10,000 1 - 3 10,037 10,101 0.69 $\%$
Barabási-Albert 10,000 2 - 3 10,042 20,105 0.72 $\%$
Barabási-Albert 10,000 5 - 3 10,043 50,103 0.77 $\%$
Erdős-Rényi 10,000 - 0.001 3 10,044 5101 0.71 $\%$
Erdős-Rényi 10,000 - 0.01 3 10,050 50,388 0.79 $\%$
Watts-Strogatz 10,000 1 0.001 3 10,054 10,132 0.84 $\%$
Watts-Strogatz 10,000 1 0.01 3 10,052 10,134 0.85 $\%$
Watts-Strogatz 10,000 2 0.001 3 10,038 20,104 0.7 $\%$
Watts-Strogatz 10,000 2 0.01 3 10,046 20,134 0.85 $\%$
Watts-Strogatz 10,000 5 0.001 3 10,039 50,090 0.63 $\%$
Watts-Strogatz 10,000 5 0.01 3 10,040 50,108 0.72 $\%$
Barabási-Albert 10,000 1 - 5 10,059 10,171 1.15 $\%$
Barabási-Albert 10,000 2 - 5 10,081 20,215 1.38 $\%$
Barabási-Albert 10,000 5 - 5 10,072 50,167 1.2 $\%$
Erdős-Rényi 10,000 - 0.001 5 10,079 5263 1.4 $\%$
Erdős-Rényi 10,000 - 0.01 5 10,067 49,998 1.08 $\%$
Watts-Strogatz 10,000 1 0.001 5 10,081 10,168 1.13 $\%$
Watts-Strogatz 10,000 1 0.01 5 10,070 10,182 1.2 $\%$
Watts-Strogatz 10,000 2 0.001 5 10,075 20,180 1.19 $\%$
Watts-Strogatz 10,000 2 0.01 5 10,071 20,190 1.24 $\%$
Watts-Strogatz 10,000 5 0.001 5 10,081 50,196 1.27 $\%$
Watts-Strogatz 10,000 5 0.01 5 10,060 50,148 1.03 $\%$
Barabási-Albert 100,000 1 - 3 100,044 100,099 0.07 $\%$
Barabási-Albert 100,000 2 - 3 100,038 200,095 0.07 $\%$
Barabási-Albert 100,000 5 - 3 100,043 500,093 0.07 $\%$
Erdős-Rényi 100,000 - 0.001 3 100,033 498,969 0.06 $\%$
Erdős-Rényi 100,000 - 0.01 3 100,044 4,998,019 0.07 $\%$
Watts-Strogatz 100,000 1 0.001 3 100,050 100,124 0.08 $\%$
Watts-Strogatz 100,000 1 0.01 3 100,045 100,096 0.07 $\%$
Watts-Strogatz 100,000 2 0.001 3 100,044 200,114 0.07 $\%$
Watts-Strogatz 100,000 2 0.01 3 100,041 200,116 0.08 $\%$
Watts-Strogatz 100,000 5 0.001 3 100,042 500,114 0.07 $\%$
Watts-Strogatz 100,000 5 0.01 3 100,045 500,116 0.08 $\%$
Barabási-Albert 100,000 1 - 5 100,063 100,167 0.11 $\%$
Barabási-Albert 100,000 2 - 5 100,069 200,175 0.12 $\%$
Barabási-Albert 100,000 5 - 5 100,065 500,159 0.12 $\%$
Erdős-Rényi 100,000 - 0.001 5 100,087 501,039 0.14 $\%$
Erdős-Rényi 100,000 - 0.01 5 100,076 4,997,365 0.13 $\%$
Watts-Strogatz 100,000 1 0.001 5 100,061 100,166 0.11 $\%$
Watts-Strogatz 100,000 1 0.01 5 100,058 100,134 0.1 $\%$
Watts-Strogatz 100,000 2 0.001 5 100,065 200,154 0.11 $\%$
Watts-Strogatz 100,000 2 0.01 5 100,059 200,154 0.11 $\%$
Watts-Strogatz 100,000 5 0.001 5 100,074 500,176 0.12 $\%$
Watts-Strogatz 100,000 5 0.01 5 100,068 500,166 0.11 $\%$
@@ 4.1.2 Open-Source Data
As post-hoc injection of specific patterns may lead to detection bias (Deprez et al., 2024) and altered topology (Egressy et al., 2024), we also include data without this injection in the experiments to evaluate GARG-AML.
We use the HI-Small and LI-Large datasets, made publicly available by Altman et al. (2023) and Egressy et al. (2024), representing a relatively small dataset with high money laundering activity and a large dataset with low money laundering activity, respectively. Table 5 summarises the number of nodes, edges and the percentage of cases labelled as money laundering. Especially LI-Large closely resembles real-world transaction networks.
@@
The two datasets are generated using the same simulator (Altman et al., 2023; Egressy et al., 2024), which models financial transactions in a virtual world involving individuals, companies, and banks. A subset of these transactions is labelled as money laundering. Since the full transaction graph is available, the authors were able to assign detailed labels, categorizing transaction streams into eight distinct types: fan-out, fan-in, gather-scatter, scatter-gather, simple cycle, random, bipartite, and stack.
@@
Table 8: Percentage ($\%$) of instances labelled as 1 in the LI-Large dataset, for the different patterns at different thresholds.
In percentage $(\%)$ 10 20 30 50 90
@@
4.2 Model Specifications
The full GARG-AML calculation pipeline consists of multiple stages, beginning with pre-processing steps that prepare the network data for analysis, followed by the computation of the GARG-AML scores. We also introduce a possible extension of the GARG-AML method that applies tree-based methods. We describe this use cases in detail below to illustrate the versatility of GARG-AML. An overview of the complete pipeline is shown in Figure 5 and summarised in Algorithm 1.
@@ 4.2 Model Specifications
1:Calculate communities using Louvain method (Blondel et al., 2008)
@@
Transaction networks often exhibit a scale-free structure (Starnini et al., 2021), which are characterized by the presence of a few nodes (accounts) with disproportionately high numbers of incoming and outgoing transactions. As a result, second-order neighbourhoods can become very large and can encompass an unmanageable number of nodes for score computation.
Therefore, we first partition the network into communities using the Louvain method (Blondel et al., 2008). This method consistently shows state-of-the-art performance for community detection, and scales well to large graphs (Lancichinetti & Fortunato, 2009; Yang et al., 2016). Communities are identified on the basis of the undirected version of the transaction network, and only intra-community edges are retained—edges between communities are discarded.
However, the Louvain method can still produce overly large communities (Fortunato & Barthélemy, 2007; Newman & Girvan, 2004). To mitigate this, we use the implementation from NetworkX (Hagberg et al., 2008), which includes a resolution parameter inspired by the improvements proposed by Traag et al. (2019). In our experiments, we set this parameter to 10, which effectively encourages the algorithm to form smaller, more manageable communities, thus maintaining tractability in the GARG-AML score computation.
We benchmark GARG-AML and its extension against established state-of-the-art methods, namely, AutoAudit (Lee et al., 2020) and FlowScope (Li et al., 2020). For both methods, we use the implementations provided by the authors and retain their default settings.
Various small adaptations are necessary. AutoAudit returns an optimal reordering of indices in the adjacency matrix. However, this reordering affects only part of the matrix and may depend on the arbitrary labelling of nodes. To ensure robustness, we repeat the experiments three times with random initialization of the node indices and report the average results. Both methods also return an ordering of transactions resulting from the $n$ different iterations. To reflect the decreasing level of certainty, the first group gets prediction $1$, the second $\frac{n-1}{n}$ and so on until the final, $n$th, group gets score $\frac{1}{n}$. All other nodes that were not returned by the methods get score 0.
@@ 4.3 Evaluation
While AUC-ROC is a widely used metric for binary classification, it may not be sufficient for anti-money laundering (AML) datasets, which are often highly imbalanced—typically with fewer than $0.1\%$ observations flagged as laundering money. To address this imbalance, we include AUC-PR, which has been shown to be a more reliable performance metric under situations of extreme label imbalance (Davis & Goadrich, 2006; Ozenne et al., 2015). In general terms, the AUC-PR is the expected average precision when a random recall threshold is selected. In other words, the AUC-PR measures the model’s ability to identify positive examples correctly while minimising false positives across all possible threshold values. The AUC-PR specifically focuses on the model’s precision in identifying the positive class at various recall thresholds.
Scalability is assessed in terms of calculation time. In addition to runtime, memory usage is also a key factor. We also report any instances where methods fail due to memory limitations. The theoretical time complexities are $O\left(k\mid E\mid\log\left(\mid V\mid\right)\right)$ for Flowscope (Li et al., 2020) and $O\left(\mid V\mid\cdot\mid E\mid+\mid V\mid^{2}\right)$ for AutoAudit (Lee et al., 2020).
For GARG-AML, the main calculations for each individual node are the selection of the second-order neighbourhood and the assignment of the neighbours to the levels (in the directed case). Assuming that we are working in sparse graphs, both steps have complexity $O\left(\mid V\mid\right)$. This brings us to a total worst case time complexity of $O\left(n^{2}\right)$. Hence, theoretically, GARG-AML scales as AutoAudit, while Flowscope is much faster. However, we can speed up the calculations by doing running them in parallel. In addition, second-order neighbours are generally much smaller than $n$, meaning that also those steps can run very quickly.
All the experiments were conducted on the Flemish Supercomputer (VSC) with an Intel Xeon Platinum 8360Y CPU and 200 GB of memory. Each experiment was constrained to a time budget of 16 hours, the maximum on the VSC.
@@ 4.4 Statistical Test
The statistical significance of the difference in ranks is evaluated using the Friedman test (Friedman, 1937, 1940) and the post hoc Nemenyi test (Nemenyi, 1963), as described by Demšar (2006). It compares the performance of the $k$ different methods across the $N$ datasets. If we take $r_{i}^{j}$ to be the rank of method $j$ on dataset $i$ for some performance metric, then the average performance across all datasets is equal to $R_{j}=\frac{1}{N}\sum_{i}r_{i}^{j}$. In our experiments, the Friedman and Nemenyi statistical tests compare the ranks of the 8 $(k)$ methods across the 66 $(N)$ synthetic datasets. We apply a significance level of $0.05$ for these tests.
@@
where critical levels $q_{\alpha}$ are provided by Demšar (2006). We visualize the results using CD diagrams (Demšar, 2006) for the AUC-ROC and AUC-PR.